What is the value of the equilibrium constant, K, for an equilibrium reaction occurring at 58.0 degrees Celsius given that the value of deltaS is 126.98 J/mol and the value of deltaH is 176.93 kJ/mol?
Answer #1

\DeltaG = \DeltaH - T\DeltaS

       = 176.93 kJ/ mol - 331*0.12698 kJ/ mol

       = -154.19 kJ/ mol

       = -154.19 *103 J/mol

\DeltaG = -RTlnK

K = -\DeltaG/ RT = 154.19 *103 J mol-1/ 331 * 8.315 JK-1mol-1

       = 154.19 *103 J mol-1 /2751.93

      = 56.03

value of equilibrium constant = 56.03